Com-Tel Inc. manufactures and assembles two models of smartphones—the Tiger Model and the Lion Model. The...

  1. Com-Tel Inc. manufactures and assembles two models of smartphones—the Tiger Model and the Lion Model. The process consists of a lean cell for each product. The data that follow concern only the Lion Model lean cell.

    For the year, Com-Tel Inc. budgeted these costs for the Lion Model production cell:

    Conversion Cost Categories Budget
    Labor $99,800
    Supplies 37,800
    Utilities 13,600
       Total $151,200

    Com-Tel plans 2,800 hours of production for the Lion Model cell for the year. The materials cost is $46 per unit. Each assembly requires 18 minutes of cell assembly time. There was no May 1 inventory for either Raw and In Process Inventory or Finished Goods Inventory.

    The following summary events took place in the Lion Model cell during May:

    1. Electronic parts were purchased to produce 9,500 Lion Model assemblies in May.
    2. Conversion costs were applied for 9,050 units of production in May.
    3. 8,870 units were completed and transferred to finished goods in May.
    4. 8,600 units were shipped to customers at a price of $216 per unit.
    If required, round your answers to the nearest cent.

    Required:

    1. Determine the budgeted cell conversion cost per hour.
    $ per hour

    2. Determine the budgeted cell conversion cost per unit.
    $ per unit

    3. Journalize the summary transactions (a) through (d). If an amount box does not require an entry, leave it blank.

    a.
    b.
    3.
    4. Sale
    4. Cost

    4. Determine the ending balance in Raw and In Process Inventory and Finished Goods Inventory.

    Raw and In Process Inventory $
    Finished Goods Inventory $
    5. Lean accounting is different from traditional accounting because it is more   and uses   control. As a result, the number of transactions are  . In many lean operations, purchased materials are charged to a  . Direct labor is  . Often, nonfinancial performance measures, such as  , are used to monitor performance.
Answer #1
1
Budgeted cell conversion cost per hour 54 per hour =151200/2800
2
Budgeted cell conversion cost per unit 16.20 per unit =54*(18/60)
3
Raw and in process inventory 437000 =9500*46
     Accounts Payable 437000
Raw and in process inventory 146610 =9050*16.20
       Conversion costs 146610
Finished goods inventory 551714 =8870*(46+16.2)
       Raw and in process inventory 551714
Accounts Receivable 1857600 =8600*216
       Sales 1857600
Cost of goods sold 534920 =8600*(46+16.2)
        Finished goods inventory 534920
4
Raw and In Process Inventory 31896 =437000+146610-551714
Finished Goods Inventory 16794 =551714-534920
5
Lean accounting is different from traditional accounting because it is more simplified and uses minimal control. As a result, the number of transactions are reduced . In many lean operations, purchased materials are charged to a “raw and in process inventory” account . Direct labor is frequently included as a conversion cost of the cell . Often, nonfinancial performance measures, such as lead time or quality measures , are used to monitor performance.